During the 2022 year there was a good rainy season and a bumper coffee bean crop yield.

The usage of the machine has improved and is estimated to continue to improve in the years to come. As such, on 31 December 2022 management have revised the remaining useful life to eight years. They have calculated the value in use to be R1 360 000. The market price is R1 850 000 with costs of R15 000.


Calculate the value of the impairment loss reversal, if any, for the year ended31 December 2022.


You can assume that the recoverable amount does not exceed what the carrying value would have been had there been no impairment loss.


1
Expert's answer
2020-08-24T14:51:45-0400

Solution:

The recoverable amount = The higher of fair value of asset less costs to sell and its value in use


"Recoverable \;Amount=R1\;850,000-R1\;15,000=R1\;835,000"

"Value\;in \;use=R1\;360,000"

"Therefore,\;the \;recoverable\;amount=R1\;835,000"


"The\; value \;of \;impairment\; loss\; reversal = Depreciated \;historical cost\; – \;Carrying \;amount"

"Depreciated\;historical \;cost
=\frac{(850,000-360,000) }{8} =61,250" annually


Historical cost for 8 yrs = "61,250\times 8=R1\;490,000"


Carrying amount = "R1\;850,000-490,000=R1\;360,000"


Therefore the value of impairment loss reversal ="R1\;490,000-R1\;360,000=R1\;130,000"